# The StudyChat Dataset: Exploring Student Dialogues With ChatGPT in an Artificial Intelligence Course

### v3

@@ 3.4 Data Statistics
Table 2: Total counts per broad dialogue act label in StudyChat. Further breakdown of specific label counts can be found in Table 3.
Label f24 s25 Total
Conceptual Questions 2204 2995 5199
Writing Request 1664 2672 4336
Provide Context 1191 1837 3028
Contextual Questions 1216 1722 2938
Verification 293 354 647
Editing 153 265 418
Off Topic 128 122 250
Misc 15 20 35
3.5 Dialogue Act Annotation
@@
In order to better understand common behavior and discourse-level trends in the conversations, we develop a DA labeling schema for student turns of the conversations. The schema is two-leveled, with 8 the broad labels and 31 specific labels, and is developed on top of existing work Wood et al. (2018) which conducted a “Wizard of Oz” experiment with programmers who were debugging alongside a simulated virtual assistant. It also mirrors our own observations after systematically reviewing the conversations, and aligns with those of Ma. et al Ma et al. (2024). The schema distinguishes the various ways students use an LLM to assist in writing aspects of their assignment and questions. We report total counts of each broad label in Table 2, total counts of specific labels in Table 3, and detail full descriptions of the schema in Table 1.
@@
To validate the accuracy of this labeling, we first sample a set of 150 utterances and develop an initial schema with labeling guidelines, including examples and descriptions for each DA. We then provide them to a research assistant, who independently labels the same set given our initial schema. We iterate on our labeling instructions and schema by computing the agreement between the two annotators and identifying sources of disagreement. After this iteration, our inter-rater agreement reached a Cohen’s kappa Cohen (1960) of 0.910 for broad level labels and 0.788 for specific labels, indicating almost perfect agreement on the broad level of DA and substantial agreement on the specific level Landis and Koch (1977). To validate this labeling, we provided these final instructions to two additional annotators who independently label the same set of utterances. We observed and average 4-way inter-rater agreement Cohen’s kappa of 0.740 for broad level agreement and 0.575 for specific level agreement, indicating substantial agreement on broad level labeling and moderate agreement on specific labeling.
@@
To apply this labeling at scale to the reminder of the dataset and since manual labeling is extraneous, we perform an LLM-prompting based approach similar to prior work Wang et al. (2023). We prompt the LLM with detailed instructions on the task, the annotation instructions, assignment instructions and the entire conversation up until the student turn to be labeled, and ask the LLM to provide the label based on our schema, along with a justification. We utilize GPT4.1 to perform this labeling, which we applied to the entire PII-filtered dataset.
@@
To validate the LLM-generated labels, we again systematically sample 154 utterances, distinct from those used in developing the initial labeling schema. We provide them to the same 4 human annotators without any further discussion or instructions. The resulting human inter-annotator agreement on the held out set is 0.578 on broad labels and 0.538 on specific labels, indicating moderate agreement between human annotators on the held out set. The average human-LLM agreement on the held out set is 0.5819 on the broad labels and 0.4927 on the specific labels, indicating moderate agreement on par with human inter-annotator agreement This drop in agreement is somewhat expected, since the new utterances in the held out set present new ambiguous, overlapping utterances to which multiple labels may apply.
4 Results and Analysis
@@
Figure 2: Average assignment scores (LLM-supported) vs. exam scores (No-LLM) for both semesters, grouped by interaction level: bottom 10% (red), middle 80% (blue), and top 10% (green). Grey dashed line shows equal outcome and red line shows data trend; top users cluster at higher scores and bottom users show wider spread and more low outliers.
4.1 Dialogue Acts as Assessment Predictors
@@
We postulate that student interactions with LLMs during assignments will have an observable influence on student course outcome. To study this hypothesis, we run a series of regression models and measure the varying effects of adding LLM usage features of increasing specificity. First, we use a single feature of the total count of utterances, then a series of features 8 corresponding to the broad DA label counts, and finally a series of 31 features corresponding to the specific DA label counts. For all models, we use a dependent variable of the score of an assessment normalized to the range $[0,1]$. In this analysis, we define assessment as one of the 7 assignments (a1-a7) or 3 course exams (e1-e3). For our baseline model, we use one independent variable, the running average of prior assessment scores in the range $[0,1]$, a well studied predictor of post assessment score hallberg2018pretest. We run each model 18 times, for 9 assessments across both semesters, excluding a1 which has no prior assessment. For the baseline model, the mean $R^{2}$ values for all regressions are $0.17343$ and $0.17281$, for the fall and spring semesters, respectively.
@@
For our first model, we add one additional independent variable to the linear regression model: the total count of all utterances inclusive of the current assignment. For exams, we sum the utterances for all assignments that were due before the exam. For these regressions, the mean $R^{2}$ was $0.1847$ for the fall and $0.1860$ for the spring, with mean $R^{2}$ improvements of $0.0064$ and $0.0178$, respectively. We run F-tests for each improvement and we find that for one assessment, spring semester a5, the addition of the usage count independent variable leads to a statically significant improvement in explained variance, with an increase of $0.0242$, $F=5.82$, $p=0.02$. In the other cases, we found no statistically significant improvement.
@@
For our next two models, we explore the influence of incorporating DA information into the regression. We expand our independent variables to include features based on the two levels of DA behavior labels, using 8 broad-labeled DAs or 31 specific-labeled DAs respectively. Specifically, we augment the baseline regression by including the counts of utterances for each of the DA labels present in the assessment target. In the same manner as above, for exams, we aggregate all assignment-specific DA counts that were due prior to the exam. With the inclusion of broad DA label counts, we observe that this inclusion has a mean $R^{2}$ of $0.2529$ for the fall and $0.2288$ for the spring semester, with a mean improvement of $0.0746$ and $0.0606$, respectively. These $R^{2}$ values indicate that DA features contribute explanatory power but F-Tests conducted for each regression do not indicate results that are statistically significant, with the highest $R^{2}$ improvement of $0.19$ and resulting F-Score of $F=1.81,p=0.09$. Repeating the same experiment with the specific DA label count features, we observe this inclusion increases the mean $R^{2}$ to $0.4868$ for the fall semester and $0.4012$ for the spring semester, with a mean $R^{2}$ increase of $0.3086$ and $0.2330$ respectively. Similar to the broad features, these $R^{2}$ values indicate that DA features contribute explanatory power but F-Tests conducted for each regression do not indicate results that are statistically significant. One assessment, a4 from the fall semester, has a borderline significant improvement of $0.4$ with an F-Score of $F=1.77,p=0.05$.
@@
Our findings in these regression analyses have mixed results in support of our hypothesis. The analysis indicates that the explanatory power of a linear regression model increases significantly when moving from a simple count of utterances to a more nuanced, qualitative representation based on DAs. However, the relative explanatory power compared to a simple, prior performance-based explainer of student outcome is only statistically significant when the utterances are represented as aggregate counts, which may simply be a confounder for student engagement or participation. This discrepancy suggests that the DA-based models are likely overfitting the data or are statistically underpowered by limited sample size. While the relationship between dialogue patterns and outcome likely exists, a more sophisticated feature selection method and modeling approach would be required to build a model that is highly predictive and robust in explaining student assessment course outcomes.
@@
We further analyze the more granular relationships between DA and assessment outcomes by examining significant coefficients in our regression models. Specifically, for the broad and specific DA regressions models, we run F-Tests to compute the significance of the coefficients for each target. Across the 36 regressions, 20 coefficients were found to be statistically significant, which we report in Table 4. For the fall semester, we observe significant coefficients only from specific-level DA regressions and only for assignment targets. Most coefficients are negatively correlated with assignment scores and largely fall into other specific DA labels, with the only positive significant correlation attributed to an off-topic DA. One possible explanation for this observation is, compared to other DAs, these labels occur relatively infrequently in the distribution, and the linear regression was likely skewed by a small subset of student behavior patterns which were uncommon.
@@
The spring semester significant coefficients are mostly distinct from the Fall, and seem to have potentially more meaningful pedagogical takeaways. One consistent theme across both semesters is that Contextual Questions acts are negatively correlated with assignment outcomes. Unique to the Spring regressions are statistically significant coefficients when the target is an exam score and the inclusion of broad-level DA features. We observe the broad feature for Conceptual Questions and Editing Request as positively correlated with e1 outcomes, indicating a learning-focus usage of LLMs being positively correlated with assessment outcome. Conversely, the Contextual Questions broad level DA was negatively correlated with Assignment 5 outcome. Since the Contextual Questions label corresponds to student questions are related to the specifics of the current conversation, this negative correlation is perhaps an indicator of student confusion. On the other hand, the Conceptual Questions label indicates a student asking a general knowledge question to the chatbot instead of something specific to the course. One possible takeaway from this finding is that educators can foster student use of LLMs for general knowledge and well-established concepts, but discourage asking for course-specific help. These suggestions may help to strike a balance for educators in fostering a modern technology-driven classroom, while preventing students from developing misconceptions from inaccurate chatbot responses.
@@
In investigating the specific DAs in the spring semester, we see Assignment 5 as a consistent target and also observe specific contextual question features as having negative correlation, with the exception of Contextual Question - Code Explanation, which has a slight positive correlation. For the spring dataset regressions, we also see different specific types of Conceptual Questions being mostly positively correlated with scores. One notable exception is that Conceptual Question - Mathematics is a strong negative indicator of a6 score. In this assignment, students were tasked with building an n-gram language model, which involved computing probability tables and deriving equations for the model. This correlation suggests that students who relied heavily on the model to understand these equations and derivations did not score well, potentially deriving incorrect solutions from the model or receiving inaccurate guidance. This observation suggests that LLM conceptual knowledge and explanations are inconsistent and that LLMs, or at least gpt4o-mini, are poor at explaining or deriving math equations. This finding suggests that educators should advise students to be cautious with LLM explanations and to inform them that while LLMs may have accurate general knowledge that the procedural derivations of these models should be carefully interpreted and validated.
4.2 Low and High Usage Users
@@
To understand how the level of user interaction influences outcome, we examine students in the bottom 10 %, middle 80% and top 10 % in terms of total utterances over the course of the semester. We will refer to these groups as low, medium and high interaction users. In the Fall dataset, the low interaction users averaged 8.6 utterances per student, and the high interaction users averaged 231.8 utterances per student. In the Spring, the low interaction averaged only 3.2 utterances per student, and the high interaction users averaged 284 utterances per student.
@@
To assess how interaction levels relate to student outcomes, we run multiple linear regression models. Identically to the previous section, we add control independent variable to each model (the running average of prior assessment scores), use a dependent variable that is the normalized score of the assessment in the range $[0,1]$, and run this configuration for our baseline regressions. For group membership, we include three additional independent binary variables which indicate whether the student was in the low, medium, or high interaction group and conduct these regressions separately for the Fall and Spring semesters.
@@
The baseline mean $R^{2}$ scores across all regressions are 0.1783 for the fall semester and 0.1682 for the spring semester. Adding the additional group features increases the mean $R^{2}$ to 0.2032 and 0.1858, respectively. This additional categorical variable yields a borderline significant improvement in explained variance for one assessment, Spring e1, $F=4.08$, $p=0.05$. On this assessment, the average scores for the low interaction, medium interaction, and high interaction users are 77.8%, 81.0%, and 88.1%, respectively. Corroborating with our finding in the previous section it appears that students with high usage, focused on conceptual questions on early assignments tend to perform better on the first exam. While the grouping-level features are statistically noteworthy, the modest gains suggest that usage patterns have limited explanatory power for overall course outcome.
@@
We compare each group by plotting individual students’ average assignment scores and average exam scores, as shown in Figure 2. We observe the minimum average assignment score (92.53), and the minimum average exam scores (72.33) for the high interaction users across both semesters are higher than the minimum of the other two groups. Moreover, the variance in the average exam and assignment scores is smaller for high interaction users. For the remaining 90% of users (low and medium interaction users), there was no substantial difference in minimum scores or variance between groups. Our analysis demonstrates that although interaction level is not a dominant predictor of outcome, high engagement with the LLM is associated with reduced variability in course outcomes and higher minimum course outcome. This pattern suggests that encouraging consistent and meaningful LLM usage, particularly among students who need additional tutoring and guidance, may help stabilize performance and raise lower-end outcomes, even if the overall effect on average scores is modest.
4.3 Clustering Students by Behavior
Figure 3: PCA of student DA features for Fall (left) and Spring (right). Points show students colored by K-means cluster. PC1 separates technical question from writing behaviors; PC2 separates context-seeking from code-oriented behaviors.

### v4

@@ 3.4. Data Statistics
3.5. Dialogue Act Annotation
@@
In order to better understand common behavior and discourse-level trends in the conversations, we develop a DA labeling schema for student turns in the conversations. The schema is two-leveled, with 8 broad labels and 31 specific labels, and is developed on top of existing work (Wood et al., 2018), which conducted a “Wizard of Oz” experiment with programmers who were debugging alongside a simulated virtual assistant. It also mirrors our own observations after systematically reviewing the conversations and aligns with those of Ma. et al (Ma et al., 2024). The schema distinguishes the various ways students use an LLM to assist in writing aspects of their assignments and ask questions. We report total counts of each broad label in Table 1.
@@
To validate the accuracy of this labeling, we first sample a set of 150 utterances and develop an initial schema with labeling guidelines, including examples and descriptions for each DA. We then provide them to a research assistant, who independently labels the same set given our initial schema. We iterate on our labeling instructions and schema by computing the agreement between the two annotators and identifying sources of disagreement. After this iteration, our inter-rater agreement reached a Cohen’s kappa (Cohen, 1960) of 0.910 for broad level labels and 0.788 for specific labels, indicating almost perfect agreement on the broad level of DA and substantial agreement on the specific level (Landis and Koch, 1977). To validate this labeling, we provided these final instructions to two additional annotators who independently label the same set of utterances. We observed an average 4-way inter-rater agreement Cohen’s kappa of 0.740 for broad level agreement and 0.575 for specific level agreement, indicating substantial agreement on broad level labeling and moderate agreement on specific labeling.
@@
To apply this labeling at scale to the remainder of the dataset and since manual labeling is extraneous, we perform an LLM-prompting-based approach similar to prior work (Wang et al., 2023; Yin et al., 2025). We prompt the LLM with detailed instructions on the task, the annotation instructions, assignment instructions, and the entire conversation up until the student turn to be labeled, and ask the LLM to provide the label based on our schema, along with a justification. We utilize GPT4.1 to perform this labeling, which we applied to the entire PII-filtered dataset.
@@
To validate the LLM-generated labels, we again systematically sample 154 utterances, distinct from those used in developing the initial labeling schema. We provide them to the same 4 human annotators without any further discussion or instructions. The resulting human inter-annotator agreement on the held out set is 0.578 on broad labels and 0.538 on specific labels, indicating moderate agreement between human annotators on the held out set. The average human-LLM agreement on the held-out set is 0.5819 on the broad labels and 0.4927 on the specific labels, indicating moderate agreement on par with human inter-annotator agreement. This drop in agreement is somewhat expected, since the new utterances in the held-out set present new ambiguous, overlapping utterances to which multiple labels may apply.
Table 2. Statistically significant regression coefficients ($p<0.05$) for predicting assessment outcomes from interaction features, grouped by semester and feature set. Only significant results are listed, and targets (assignments and exams) abbreviated.
Semester Feature Set Feature Name Target Coef. $p$ $n$
F24 Broad No significant coefficients
Specific Contextual Questions: Code Explanation A3 0.0065-0.0065 $0.021$ $72$
Contextual Questions: Other A3 0.0156-0.0156 $0.034$ $72$
Off Topic: Chit-Chat A3 $0.0235$ $0.043$ $72$
Provide Context: Code A4 0.0426-0.0426 $0.001$ $66$
Verification: Other A4 0.0291-0.0291 $0.007$ $66$
Conceptual Questions: Other Concept A5 0.0575-0.0575 $0.025$ $64$
Writing Req: Other A5 0.0590-0.0590 $0.033$ $64$
S25 Broad Conceptual Questions E1 $0.0133$ $0.024$ $104$
Editing Req E1 $0.0296$ $0.017$ $104$
Contextual Questions A5 0.0020-0.0020 $0.005$ $72$
Specific Conceptual Questions: Python Library E1 $0.0234$ $0.010$ $104$
Contextual Questions: Assignment Clarif. A3 0.0034-0.0034 $0.021$ $103$
Conceptual Questions: Other Concept A3 $0.0095$ $0.029$ $103$
Writing Req: Conversion E2 $0.0028$ $0.045$ $119$
Contextual Questions: Assignment Clarif. A5 0.0072-0.0072 $0.039$ $72$
Contextual Questions: Code Explanation A5 $0.0009$ $0.044$ $72$
Contextual Questions: Other A5 0.0075-0.0075 $0.017$ $72$
Conceptual Questions: Other Concept A5 $0.0016$ $0.019$ $72$
Conceptual Questions: Mathematics A6 0.1062-0.1062 $0.002$ $74$
Off Topic: Gratitude E3 0.0337-0.0337 $0.025$ $119$
4. Results and Analysis
@@
4.1. Dialogue Acts as Assessment Predictors
@@
We postulate that student interactions with LLMs during assignments will have an observable influence on student course outcomes. To study this hypothesis, we run a series of regression models and measure the varying effects of adding LLM usage features of increasing specificity. First, we use a single feature of the total count of utterances, then a series of 8 features corresponding to the broad DA label counts, and finally a series of 31 features corresponding to the specific DA label counts. For all models, we use a dependent variable of the score of an assessment normalized to the range $[0,1]$. In this analysis, we define assessment as one of the 7 assignments (a1-a7) or 3 course exams (e1-e3). For our baseline model, we use one independent variable, the running average of prior assessment scores in the range $[0,1]$, a well-studied predictor of post assessment score (Hallberg et al., 2018). We run each model 18 times, for 9 assessments across both semesters, excluding a1, which has no prior assessment. For the baseline model, the mean $R^{2}$ values for all regressions are $0.17343$ and $0.17281$, for the fall and spring semesters, respectively.
@@
For our first model, we add one additional independent variable to the linear regression model: the total count of all utterances, inclusive of the current assignment. For exams, we sum the utterances for all assignments that were due before the exam. For these regressions, the mean $R^{2}$ was $0.1847$ for the fall and $0.1860$ for the spring, with mean $R^{2}$ improvements of $0.0064$ and $0.0178$, respectively. We run F-tests for each improvement, and we find that for one assessment, spring semester a5, the addition of the usage count independent variable leads to a statistically significant improvement in explained variance, with an increase of $0.0242$, $F=5.82$, $p=0.02$. In the other cases, we found no statistically significant improvement.
@@
For our next two models, we explore the influence of incorporating DA information into the regression. We expand our independent variables to include features based on the two levels of DA behavior labels, using 8 broad-labeled DAs or 31 specific-labeled DAs, respectively. Specifically, we augment the baseline regression by including the counts of utterances for each of the DA labels present in the assessment target. In the same manner as above, for exams, we aggregate all assignment-specific DA counts that were due prior to the exam. With the inclusion of broad DA label counts, we observe that this inclusion has a mean $R^{2}$ of $0.2529$ for the fall and $0.2288$ for the spring semester, with a mean improvement of $0.0746$ and $0.0606$, respectively. These $R^{2}$ values indicate that DA features contribute explanatory power, but F-Tests conducted for each regression do not indicate results that are statistically significant, with the highest $R^{2}$ improvement of $0.19$ and resulting F-Score of $F=1.81,p=0.09$. Repeating the same experiment with the specific DA label count features, we observe that this inclusion increases the mean $R^{2}$ to $0.4868$ for the fall semester and $0.4012$ for the spring semester, with a mean $R^{2}$ increase of $0.3086$ and $0.2330$, respectively. Similar to the broad features, these $R^{2}$ values indicate that DA features contribute explanatory power, but F-Tests conducted for each regression do not indicate results that are statistically significant. One assessment, a4 from the fall semester, has a borderline significant improvement of $0.4$ with an F-Score of $F=1.77,p=0.05$.
@@
Our findings in these regression analyses have mixed results in support of our hypothesis. The analysis indicates that the explanatory power of a linear regression model increases significantly when moving from a simple count of utterances to a more nuanced, qualitative representation based on DAs. However, the relative explanatory power compared to a simple, prior performance-based explainer of student outcome is only statistically significant when the utterances are represented as aggregate counts, which may simply be a confounder for student engagement or participation. This discrepancy suggests that the DA-based models are likely overfitting the data or are statistically underpowered by a limited sample size. While the relationship between dialogue patterns and outcome likely exists, a more sophisticated feature selection method and modeling approach would be required to build a model that is highly predictive and robust in explaining student assessment course outcomes.
@@
We further analyze the more granular relationships between DA and assessment outcomes by examining significant coefficients in our regression models. Specifically, for the broad and specific DA regression models, we run F-Tests to compute the significance of the coefficients for each target. Across the 36 regressions, 20 coefficients were found to be statistically significant, which we report in Table 2. For the fall semester, we observe significant coefficients only from specific-level DA regressions and only for assignment targets. Most coefficients are negatively correlated with assignment scores and largely fall into other specific DA labels, with the only positive significant correlation attributed to an off-topic DA. One possible explanation for this observation is that, compared to other DAs, these labels occur relatively infrequently in the distribution, and the linear regression was likely skewed by a small subset of student behavior patterns that were uncommon.
@@
The spring semester’s significant coefficients are mostly distinct from the Fall, and seem to have potentially more meaningful pedagogical takeaways. One consistent theme across both semesters is that Contextual Questions acts are negatively correlated with assignment outcomes. Unique to the Spring regressions are statistically significant coefficients when the target is an exam score and the inclusion of broad-level DA features. We observe the broad feature for Conceptual Questions and Editing Request as positively correlated with e1 outcomes, indicating a learning-focused usage of LLMs being positively correlated with assessment outcomes. Conversely, the Contextual Questions broad level DA was negatively correlated with Assignment 5 outcome. Since the Contextual Questions label corresponds to student questions that are related to the specifics of the current conversation, this negative correlation is perhaps an indicator of student confusion. On the other hand, the Conceptual Questions label indicates a student asking a general knowledge question to the chatbot instead of something specific to the course. One possible takeaway from this finding is that educators can foster student use of LLMs for general knowledge and well-established concepts, but discourage asking for course-specific help. These suggestions may help to strike a balance for educators in fostering a modern technology-driven classroom, while preventing students from developing misconceptions from inaccurate chatbot responses.
@@
In investigating the specific DAs in the spring semester, we see Assignment 5 as a consistent target and also observe specific contextual question features as having negative correlation, with the exception of Contextual Question - Code Explanation, which has a slight positive correlation. For the spring dataset regressions, we also see different specific types of Conceptual Questions being mostly positively correlated with scores. One notable exception is that Conceptual Question - Mathematics is a strong negative indicator of a6 score. In this assignment, students were tasked with building an $n$-gram language model, which involved computing probability tables and deriving equations for the model. This correlation suggests that students who relied heavily on the model to understand these equations and derivations did not score well, potentially deriving incorrect solutions from the model or receiving inaccurate guidance. This observation suggests that LLM conceptual knowledge and explanations are inconsistent and that LLMs, or at least gpt4o-mini, are poor at explaining or deriving math equations. This finding suggests that educators should advise students to be cautious with LLM explanations and to inform them that while LLMs may have accurate general knowledge, the procedural derivations of these models should be carefully validated.
Figure 3. Average assignment scores (LLM-supported) vs. exam scores (No-LLM) for Fall 2024 (left, circles) and Spring 2025 (right, squares), grouped by interaction level: bottom 10% (red), middle 80% (blue), and top 10% (green). Grey dashed line shows equal outcome, and red line shows the data trend for each semester; top users cluster at higher scores, and bottom users show a wider spread and more low outliers across both semesters.
4.2. Low and High Usage Users
@@
To understand how the level of user interaction influences outcome, we examine students in the bottom 10%, middle 80%, and top 10% in terms of total utterances over the course of the semester. We will refer to these groups as low, medium, and high interaction users. In the Fall dataset, the low interaction users averaged 8.6 utterances per student, and the high interaction users averaged 231.8 utterances per student. In the Spring, the low interaction averaged only 3.2 utterances per student, and the high interaction users averaged 284 utterances per student.
@@
To assess how interaction levels relate to student outcomes, we run multiple linear regression models. Identically to the previous section, we add a control independent variable to each model (the running average of prior assessment scores), use a dependent variable that is the normalized score of the assessment in the range $[0,1]$, and run this configuration for our baseline regressions. For group membership, we include three additional independent binary variables that indicate whether the student was in the low, medium, or high interaction group and conduct these regressions separately for the Fall and Spring semesters.
@@
The baseline mean $R^{2}$ scores across all regressions are 0.1783 for the fall semester and 0.1682 for the spring semester. Adding the additional group features increases the mean $R^{2}$ to 0.2032 and 0.1858, respectively. This additional categorical variable yields a borderline significant improvement in explained variance for one assessment, Spring e1, $F=4.08$, $p=0.05$. On this assessment, the average scores for the low interaction, medium interaction, and high interaction users are 77.8%, 81.0%, and 88.1%, respectively. Corroborating with our finding in the previous section, it appears that students with high usage, focused on conceptual questions on early assignments, tend to perform better on the first exam. While the grouping-level features are statistically noteworthy, the modest gains suggest that usage patterns have limited explanatory power for the overall course outcome.
@@
We compare each group by plotting individual students’ average assignment scores and average exam scores, as shown in Figure 3. We observe the minimum average assignment score (92.53), and the minimum average exam scores (72.33) for the high interaction users across both semesters are higher than the minimum of the other two groups. Moreover, the variance in the average exam and assignment scores is smaller for high interaction users. For the remaining 90% of users (low and medium interaction users), there was no substantial difference in minimum scores or variance between groups. Our analysis demonstrates that although interaction level is not a dominant predictor of outcome, high engagement with the LLM is associated with reduced variability in course outcomes and higher minimum course outcomes. This pattern suggests that encouraging consistent and meaningful LLM usage, particularly among students who need additional tutoring and guidance, may help stabilize performance and raise lower-end outcomes, even if the overall effect on average scores is modest.
4.3. Clustering Students by Behavior
Fall 2024 Spring 2025
Specific Dialogue Act C0 C1 C2 C3 C0 C1 C2 C3
Conceptual Questions: Python Library $4.2$ $26.3$ $16.0$ $2.4$ $1.2$ $3.0$ $15.7$ $0.0$
Writing Request: Write Code $24.4$ $11.3$ $9.3$ $17.8$ $7.9$ $22.3$ $12.1$ $0.0$
Writing Request: Write English $8.6$ $1.7$ $4.0$ $16.8$ $5.8$ $13.0$ $5.1$ $0.0$
Conceptual Questions: Programming Language $4.4$ $15.6$ $6.5$ $1.5$ $0.9$ $1.4$ $10.2$ $0.0$
Writing Request: Code/Data Conversion $2.6$ $2.1$ $4.0$ $14.5$ $0.6$ $7.1$ $3.4$ $0.0$
Conceptual Questions: Computer Science $4.1$ $13.1$ $7.7$ $8.1$ $2.6$ $4.3$ $11.3$ $0.0$
Contextual Questions: Code Explanation $3.7$ $4.0$ $9.4$ $2.7$ $4.3$ $4.6$ $6.3$ $3.0$
Contextual Questions: Assignment Clarification $5.9$ $2.7$ $8.8$ $3.2$ $4.1$ $10.3$ $6.3$ $3.0$
Provide Context: Error Message $8.0$ $3.2$ $2.9$ $1.8$ $8.5$ $5.0$ $3.5$ $0.0$
Provide Context: Other $7.4$ $2.3$ $3.9$ $3.7$ $23.9$ $4.8$ $3.1$ $0.0$
Cluster Size 25 16 27 16 5 68 43 3
Figure 4. PCA of DA features for Fall (top-left) and Spring (top-right). The table (bottom) summarizes the cluster distributions and shows the percentage of specific dialogue acts for the top 10 most common specific DA across all clusters. We see the type of question and writing requests as consistent distinguishing features across clusters. See appendix for further cluster details.